OpenStudy (anonymous):

How do I solve log9(I can't find my arrow that points down, but the base 9 is smaller at the bottom. How do I solve log9 81 = x

OpenStudy (shubhamsrg):

your eqn indirectly states 9^x = 81 so,,is it easier now ?

OpenStudy (anonymous):

Yes! Thanks! So to solve it, I find what x equal, right?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

9^x =9^2 so than x= ?

OpenStudy (anonymous):

Thank you.(:

jhonyy9 (jhonyy9):

yw good luck bye